### v1

# LAPO: Internalizing Reasoning Efficiency via Length-Adaptive Policy Optimization

Abstract
Large reasoning models have achieved remarkable performance through extended chain-of-thought sequences, yet this computational freedom leads to excessive token generation even for simple problems. We present Length-Adaptive Policy Optimization (LAPO), a novel framework that transforms reasoning length control from an external constraint into an intrinsic model capability. Unlike existing approaches that impose rigid limits or rely on post-hoc interventions, LAPO enables models to internalize an understanding of appropriate reasoning depth through a two-stage reinforcement learning process. In the first stage, models learn natural reasoning patterns by discovering the statistical distribution of successful solution lengths. The second stage leverages these patterns as meta-cognitive guidance, embedding them directly within the model’s reasoning context to ensure inference-time flexibility. Experiments on mathematical reasoning benchmarks demonstrate that LAPO reduces token usage by up to 40.9% while improving accuracy by 2.3%. Our analysis reveals that models trained with LAPO develop emergent abilities to allocate computational resources based on problem complexity, achieving efficient reasoning without sacrificing quality.
1 Introduction
Recent advances in large reasoning models have demonstrated remarkable capabilities through extended chain-of-thoughts (Wei et al., 2022; Jaech et al., 2024; Guo et al., 2025). However, this computational freedom comes with a critical limitation: these models often generate excessively verbose reasoning chains regardless of problem complexity, leading to significant computational overhead, a phenomenon known as ”overthinking” (Min et al., 2024). While beneficial for complex problems, this verbosity introduces substantial inefficiencies, making practical deployment challenging.
Existing approaches to address this challenge fall into three main categories, each with inherent limitations. Direct length reduction methods either rely on reward design (Yang et al., 2025; Huang et al., 2025) that can cause over-shortening and accuracy degradation, or impose hard length constraints (Aggarwal & Welleck, 2025; Hou et al., 2025) that lack adaptability across problem types. Dynamic early-stopping approaches (Qiao et al., 2025; Muennighoff et al., 2025) make real-time termination decisions but often truncate mid-reasoning, disrupting the thinking process. Adaptive thinking methods (Lou et al., 2025; Zhang et al., 2025; Fang et al., 2025) enable models to switch between thinking and non-thinking modes but operate at a coarse granularity.
The fundamental limitation of these approaches lies in their treatment of length control as an external constraint imposed upon the reasoning process. This paradigm inherently conflicts with the nature of mathematical reasoning, where each problem possesses its own intrinsic complexity that naturally determines the required reasoning depth. Current methods fail to recognize that when models successfully solve problems, they naturally converge to certain reasoning lengths that reflect this intrinsic complexity. The challenge is not to impose arbitrary limits, but to help models discover and internalize these natural reasoning patterns.
We propose a paradigm shift: instead of constraining reasoning through external mechanisms, we enable models to learn from their own successful reasoning patterns and develop an internal sense of appropriate reasoning depth. Our key insight is that the distribution of reasoning lengths in correct solutions contains valuable information about how much thinking each problem genuinely requires. By capturing these patterns during training and teaching models to anticipate the appropriate reasoning budget before they begin solving, we can transform length control from an external limitation into an intrinsic capability.
Figure 1: Overview of Length-Adaptive Policy Optimization (LAPO) and its superior performance. The LAPO framework (left) trains a model in two stages: first discovering natural reasoning lengths, then internalizing them as self-proposed budgets. This process enables our models (LAPO-I) to achieve a state-of-the-art balance between accuracy and efficiency (right), surpassing existing methods by operating in the desirable top-left region of the performance plot.
We introduce Length-Adaptive Policy Optimization (LAPO), a two-stage reinforcement learning framework that progressively builds this adaptive capability. In the first stage, we design length-aware rewards that encourage efficiency while maintaining accuracy. During this process, we collect statistical patterns from successful solutions, specifically focusing on the reasonable length range where most correct answers naturally fall. This reveals the inherent reasoning requirements of each problem without imposing artificial constraints. In the second stage, we leverage these discovered patterns to provide explicit guidance to the model. By incorporating target length information directly into the problem prompt, we enable models to plan their reasoning trajectory before beginning the solution process. Crucially, to ensure models can reason adaptively without requiring predefined lengths at inference time, we embed the length constraint as a self-declarative statement immediately after the ¡think¿ token. This technique reframes the budget not as an external command, but as part of the model’s own internal reasoning plan. By learning to generate a solution that aligns with its self-proposed budget, the model is incentivized to internalize the link between problem complexity and resource allocation, allowing it to reason flexibly and efficiently when deployed.
LAPO fundamentally differs from existing approaches by recognizing that efficient reasoning requires understanding problem-specific computational needs rather than following rigid rules. Our two-stage design enables a natural progression: models first learn what constitutes appropriate reasoning depth through experience, then develop the ability to anticipate these requirements proactively. This approach mirrors how human experts develop intuition about problem complexity, allocating mental effort proportionally to task demands.
Extensive experiments validate the effectiveness of our approach. LAPO achieves remarkable efficiency gains, reducing token usage by up to 40.9% while simultaneously improving accuracy by 2.3% on mathematical reasoning benchmarks (see Figure 1). Our analysis reveals that this improvement stems from the model’s ability to distinguish between problems requiring elaborate derivations versus those needing only brief calculations. The training dynamics demonstrate smooth convergence in both stages, with models maintaining stable accuracy even as they learn increasingly precise length control. These results confirm that when models learn from their own successful patterns rather than arbitrary constraints, they develop more robust and efficient reasoning strategies.
Our main contributions are:
  • We propose LAPO, a novel two-stage reinforcement learning framework that transforms length control from an external constraint into an intrinsic reasoning capability, enabling models to adaptively allocate computational resources based on problem complexity.
  • We introduce a training methodology that combines statistical analysis of successful reasoning patterns with contextual length guidance embedded within the reasoning process, allowing models to internalize length-adaptive behaviors while maintaining inference-time flexibility.
  • We demonstrate through extensive experiments that LAPO achieves substantial efficiency gains (up to 40.9% token reduction) while improving accuracy, revealing that models can develop emergent meta-cognitive abilities for reasoning budget allocation.
2 Related Works
2.1 Test-Time Scaling in Large Language Models
Increasing test-time computation has consistently been shown to improve performance in complex reasoning tasks, mathematical problem-solving, and code generation (Wu et al., 2024; Wang et al., 2022; Wei et al., 2022; Guo et al., 2025). Test-time scaling laws indicate predictable performance gains from increasing inference computation, either by generating more reasoning chains or longer ones (Wu et al., 2024; Snell et al., 2024; Jaech et al., 2024). Prominent approaches include parallel sampling of multiple reasoning paths (Wang et al., 2022), tree-based search (Yao et al., 2023; Wu et al., 2024), and iterative refinement techniques (Snell et al., 2024; Welleck et al., 2024).
Recent reasoning models such as OpenAI’s O1 and DeepSeek’s R1-style models (Jaech et al., 2024; Guo et al., 2025) simplify test-time scaling by generating extended reasoning traces through reinforcement learning with verifiable rewards (RLVR), encouraging deep thinking behaviors such as broad exploration and feasibility checks (Gandhi et al., 2025). However, these extended reasoning behaviors often lead to much longer reasoning traces, sometimes several times longer than those produced by short CoT models (Sui et al., 2025; Chen et al., 2024), creating an “overthinking” issue that largely increases inference costs (Kumar et al., 2025). Several recent works have shown that this extended reasoning often includes redundant or unnecessary verification and reflection, even on simple problems (Wang et al., 2025). Despite their promising results, existing methods lack precise and dynamic control over the length of generated reasoning chains, resulting in often suboptimal performance or unrealized potential efficiency gains.
2.2 Efficient Long Chain-of-Thought LLM
While test-time scaling with long CoT significantly improves accuracy, it comes at the cost of computational inefficiency. In particular, reasoning models often produce verbose and unnecessary reasoning when solving simple problems—a phenomenon commonly referred to as overthinking (Sui et al., 2025). To address the overthinking phenomenon in reasoning models, various methods have been proposed following three main strategies. Prompt-based methods attempt to control response length by incorporating instructions directly into prompts (Xu et al., 2025a), but cannot achieve precise length control. Training-based methods include supervised fine-tuning approaches that collect datasets with variable lengths (Han et al., 2024; Kang et al., 2025; Ma et al., 2025; Xia et al., 2025) and RL-based methods that incorporate length penalties into reward functions (Muennighoff et al., 2025; Yeo et al., 2025; Luo et al., 2025a; Xu et al., 2025b). However, these methods fail to control length according to users’ requirements or problem complexity. Router-based methods train separate classifiers to route queries between fast and reasoning models (Chuang et al., 2024; Ong et al., 2024), but require additional computational overhead. Recent advances in token budget control have introduced more sophisticated approaches. Works like L1 (Aggarwal & Welleck, 2025) and Elastic Reasoning (Xu et al., 2025b) can more precisely control output length under given token budgets, yet they fail to enable autonomous estimation of appropriate response lengths for different problems.
In contrast to these prior approaches, our LAPO framework uniquely combines autonomous budget estimation and precise length control capabilities through a two-stage reinforcement learning design. Unlike existing methods that rely on external truncation mechanisms or require manual budget specification, LAPO trains models to intrinsically learn appropriate reasoning lengths while maintaining reasoning completeness and logical coherence. This endogenous length control capability enables problem-adaptive token budget allocation, achieving significant efficiency improvements while maintaining or enhancing reasoning performance.
3 Method
We present Length-Adaptive Policy Optimization (LAPO), a framework that enables reasoning models to internalize efficient reasoning as an intrinsic capability. Our approach fundamentally differs from existing methods by teaching models to develop an internal understanding of appropriate reasoning depth, rather than imposing external constraints. We achieve this through a carefully designed two-stage training process that first discovers natural reasoning patterns, then transforms these patterns into an internalized capability.
3.1 Overview
Consider a reasoning model generating response $r$ for problem $q$. While current models produce high-quality solutions, they lack awareness of computational efficiency, often generating responses far exceeding necessary length. Our goal is to train models that autonomously determine appropriate reasoning lengths while maintaining solution quality.
Our key insight is that successful problem solutions naturally exhibit certain length distributions that reflect intrinsic problem complexity. Rather than viewing these patterns as constraints to enforce, we treat them as signals that teach models about reasoning depth requirements. LAPO employs a two-stage approach illustrated in Figure 2: the Discovery stage explores natural reasoning patterns through length-aware rewards, while the Internalization stage transforms these patterns into adaptive reasoning behavior.
Figure 2: The LAPO framework consists of two stages: (1) Discovery stage learns natural reasoning patterns by rewarding efficient correct solutions and collecting length statistics; (2) Internalization stage embeds these statistics as self-proposed plans within the model’s reasoning context, teaching models to internalize efficient reasoning.
3.2 Discovery Stage: Learning Natural Reasoning Patterns
The Discovery stage aims to uncover inherent relationships between problems and their natural reasoning lengths through GRPO training with a carefully designed reward mechanism that encourages efficient exploration while maintaining correctness.
Extracting Statistics from GRPO Rollouts.
During GRPO training, we generate $N$ rollout responses for each problem $q$ in the training batch. From these rollouts, we collect the lengths of responses that produce correct answers:
$\mathcal{L}_{q}=\{|r_{i}|:\mathbb{I}(y_{i}=y_{\text{gold}})=1,i\in[1,N]\}$ (1)
where $y_{i}$ is the predicted answer from the $i$-th rollout response $r_{i}$. This collection, extracted directly from the GRPO sampling process, represents natural variation in successful reasoning lengths.
We derive two key statistics from these rollouts. First, we establish a reasonable length range using percentiles to filter outliers while preserving central tendencies:
$[L_{\min},L_{\max}]=[\text{Percentile}_{30}(\mathcal{L}_{q}),\text{Percentile}_{70}(\mathcal{L}_{q})]$ (2)
Second, we create a problem-to-length mapping that will guide the Internalization stage:
$\mathcal{M}:q\mapsto L_{\text{median}}(q)=\text{Median}(\mathcal{L}_{q})$ (3)
For problems without correct solutions in the current rollouts, we set $\mathcal{M}(q)=4096$ (maximum sequence length) to encourage comprehensive exploration in subsequent episodes.
Length-Aware Reward Design.
We employ a composite reward function balancing accuracy and efficiency:
$R_{1}(r_{i},q)=\mathbb{I}(y_{i}=y_{\text{gold}})+\alpha\cdot R_{\text{length-1}}(r_{i},q)$ (4)
The length component operates on a crucial principle—only correct responses receive length-based rewards:
$R_{\text{length-1}}(r_{i},q)=\begin{cases}1.0&\text{if correct and }L_{\min}\leq|r_{i}|\leq L_{\max},\\ \max(0,1-d_{i}/100)&\text{if correct and }|r_{i}|\notin[L_{\min},L_{\max}],\\ 0&\text{if incorrect;}\end{cases}$ (5)
where $d_{i}=\min(||r_{i}|-L_{\min}|,||r_{i}|-L_{\max}|)$. This design creates gradients guiding models toward efficient lengths while allowing flexibility for complex problems. Throughout the Discovery stage, we continuously update $\mathcal{M}$ after each GRPO training step to reflect evolving model capabilities.
3.3 Internalization Stage: Length-Aware Efficient Reasoning
The Internalization stage transforms discovered patterns into internalized capabilities through continued GRPO training with modified prompts and rewards.
Length-Conditioned Rollout.
We augment each problem prompt with explicit length guidance:
$\text{prompt}_{q}^{\prime}=\text{prompt}_{q}+\text{``{<think>} I will answer the question with }n\text{ tokens.''}$.” (6)
where $n=\mathcal{M}(q)$ from the Discovery stage. This embeds length awareness within the reasoning context, helping models perceive computational budgets as intrinsic to thinking rather than external constraints.
Length-Adherence Reward.
To encourage the model to follow its self-declared reasoning budget from Equation 6, the Internalization stage employs a precision-focused reward function. This function is designed to reward the alignment between the model’s output length and its self-declared budget n. The total reward is defined as:
$R_{2}(r_{i},q^{\prime})=\mathbb{I}(y_{i}=y_{\text{gold}})+\beta\cdot R_{\text{Length-2}}(r_{i},q^{\prime})$ (7)
where the adherence component, $R_{\text{Length-2}}$, is only granted for correct solutions:
$R_{\text{Length-2}}(r_{i},n)=\begin{cases}\exp\left(-\frac{(|r_{i}|-n)^{2}}{2\sigma^{2}}\right)&\text{if correct},\\ 0&\text{otherwise};\end{cases}$ (8)
This Gaussian-inspired reward positively reinforces solutions that are both correct and consistent with the intended reasoning depth. By rewarding adherence to the self-proposed plan, this mechanism guides the model to internalize the relationship between problem complexity and an appropriate computational budget, rather than merely tracking an external signal.
Algorithm 1 Length-Adaptive Policy Optimization (LAPO)
1:Input: Base model $\pi_{\theta}$, training data $\mathcal{D}$, hyperparameters $\alpha,\beta,\sigma,E_{1},E_{2}$
2:Output: Length-adaptive model $\pi_{\theta}^{*}$
3:
4:// Discovery Stage
5:for episode $e=1$ to $E_{1}$ do
6:     Sample batch $\mathcal{B}\subset\mathcal{D}$
7:     for each problem $q\in\mathcal{B}$ do
8:         Generate $N$ rollouts: $\{r_{1},...,r_{N}\}\sim\pi_{\theta}(q)$
9:         Collect correct lengths: $\mathcal{L}_{q}=\{|r_{i}|:y_{i}=y_{\text{gold}}\}$
10:         Compute range: $[L_{\min},L_{\max}]=[\text{P}_{30}(\mathcal{L}_{q}),\text{P}_{70}(\mathcal{L}_{q})]$
11:         Update mapping: $\mathcal{M}(q)=\text{Median}(\mathcal{L}_{q})$
12:         Compute rewards: $R_{1}(r_{i},q)=\mathbb{I}(y_{i}=y_{\text{gold}})+\alpha\cdot R_{\text{length-1}}(r_{i},q)$
13:     end for
14:     Update $\pi_{\theta}$ using GRPO with rewards $R_{1}$
15:end for
16:
17:// Internalization Stage
18:for episode $e=1$ to $E_{2}$ do
19:     Sample batch $\mathcal{B}\subset\mathcal{D}$
20:     for each problem $q\in\mathcal{B}$ do
21:         Augment prompt: $q^{\prime}=q+\text{``{<think>} I will answer the question with }\mathcal{M}(q)\text{ tokens.''}$.”
22:         Generate $N$ rollouts: $\{r_{1},...,r_{N}\}\sim\pi_{\theta}(q^{\prime})$
23:         Compute rewards: $R_{2}(r_{i},q^{\prime})=\mathbb{I}(y_{i}=y_{\text{gold}})+\beta\cdot R_{\text{Length-2}}(r_{i},q^{\prime})$
24:         Update mapping $\mathcal{M}(q)$ using dual-strategy (Eq. 9)
25:     end for
26:     Update $\pi_{\theta}$ using GRPO with rewards $R_{2}$
27:end for
28:return $\pi_{\theta}^{*}$
Internalization via In-Context Guidance.
A cornerstone of our framework is how it fosters genuine internalization, enabling inference-time flexibility without explicit length targets. The key lies in the design of the augmented prompt (Equation 6). By placing the length guidance ”I will answer the question with $n$ tokens.” immediately after the <think> token, we transform an external constraint into an intrinsic part of the model’s cognitive process.
During the Internalization stage, we refine $\mathcal{M}$ based on new GRPO rollouts with a dual-strategy update:
$\mathcal{M}(q)=\begin{cases}\text{Median}(\mathcal{L}_{q}^{(t)})&\text{if previously unsolved}\\ \min(\mathcal{M}(q),\text{Median}(\mathcal{L}_{q}^{(t)}))&\text{if previously solved}\end{cases}$ (9)
This ensures newly solved problems establish reasonable benchmarks while previously solved problems gravitate toward more efficient solutions.
3.4 Training Pipeline
We present the complete LAPO training procedure in Algorithm 1. LAPO employs GRPO across both stages with the following pipeline:
Discovery Stage
(Lines 4-13): The model explores natural reasoning patterns through GRPO training with length-aware rewards. For each problem in the training batch, we generate multiple rollouts and extract statistics from successful responses. The mapping $\mathcal{M}$ is continuously updated to capture the evolving understanding of appropriate reasoning lengths. This stage runs for $E_{1}$ epochs, allowing the model to discover problem-specific length patterns through self-supervised exploration.
Internalization Stage
(Lines 15-24): The model learns to internalize efficient reasoning by incorporating discovered length patterns into the training process. Each problem prompt is augmented with target length information derived from the Discovery stage. The placement of this guidance within the <think> block encourages the model to treat the budget as part of its own reasoning plan, which fosters genuine length awareness rather than rote instruction following. The dual-strategy update mechanism refines the mapping $\mathcal{M}$ throughout training, allowing newly solved problems to establish benchmarks while encouraging efficiency improvements for previously solved ones.
This progressive design mirrors cognitive development: first gaining experience about appropriate reasoning depth through practice, then learning to anticipate these requirements proactively. The embedding of guidance as a self-declared plan is the key mechanism that bridges this gap from experience to proactive anticipation. By making efficiency an intrinsic part of reasoning, LAPO creates models that naturally adapt computational investment to match problem demands.
4 Experiment Setup
Training Details.
We train our models on a mixed dataset of 10,000 mathematical problems. This dataset comprises 6,000 examples randomly sampled from the DeepScaleR-Preview-Dataset (Luo et al., 2025b), which sources problems from competitive math contests (e.g., AIME, AMC) and benchmarks (e.g., Omni-Math (Gao et al., 2024), STILL (Min et al., 2024)). The remaining 4,000 examples are drawn from the intermediate levels (3-5) of the MATH dataset (Hendrycks et al., 2020) to ensure a balanced difficulty distribution. For the base models, we experiment with DeepSeek-R1-1.5B(Guo et al., 2025), a strong instruction-tuned reasoner, and DeepScaleR-1.5B-Preview(Luo et al., 2025b), an advanced RL-finetuned version known for high-quality but verbose outputs.
All models are trained using the Group Relative Policy Optimization (GRPO) algorithm. The LAPO framework is applied in two sequential stages. Stage 1 (LAPO-Discovery) is trained for 240 steps (approx. 3 epochs) with the length-aware reward $R_{1}$ (Eq. 4), where hyperparameter $\alpha$ is 0.7. The resulting model then serves as the initialization for Stage 2 (LAPO-Internalization), which is trained for another 240 steps using the plan-adherence reward $R_{2}$ (Eq. 7) with $\beta$ set to 0.8. Due to computational constraints, the maximum context length is limited to 4,096 tokens during training. More training details are provided in Appendix A.1.
Evaluation Details.
We evaluate the trained models on four challenging benchmarks representing a spectrum of mathematical difficulty: MATH-500 (Hendrycks et al., 2020), AIME2024, AMC23, and Olympiad-Bench (He et al., 2024). Following standard practices (Guo et al., 2025), we use a sampling temperature of 0.6, a top-p of 0.95, and a maximum generation length of 32,768 tokens. For each benchmark, we report Pass@1 accuracy and the average number of tokens per response. The number of samples per query is 4 for MATH-500 and Olympiad-Bench, and 32 for the more competitive AIME2024 and AMC23.
Baselines.
We benchmark LAPO against two classes of baselines: the foundational models we build upon and existing methods designed for efficient reasoning. First, we evaluate the Base Models to establish a performance starting point. These include DeepSeek-R1-1.5B(Guo et al., 2025), a strong instruction-tuned model, and DeepScaleR-1.5B-Preview(Luo et al., 2025b), a powerful RL-finetuned reasoner known for its high-quality but often verbose outputs. Both are evaluated without any length-control modifications. Second, we compare against state-of-the-art Efficient Reasoning Baselines that represent two distinct technical approaches. For methods focusing on direct length budgeting, we evaluate L1-Exact and L1-Max(Aggarwal & Welleck, 2025). Both train a model to adhere to a user-specified length constraint by penalizing deviations from the target budget, with L1-Exact enforcing a precise match and L1-Max treating the budget as a soft upper limit. For the iterative pruning strategy, we compare against ThinkPrune(Hou et al., 2025). This method trains a model to shorten its reasoning by imposing a hard token limit during its iterative RL process, where any response exceeding the limit receives a zero reward. We evaluate its two publicly available variants, ThinkPrune-iter2k and ThinkPrune-4k.
5 Results and Analysis
We present comprehensive experimental results and analysis to validate LAPO’s effectiveness and understand its underlying mechanisms. We begin with the main results (Section 5.1), benchmarking LAPO against baselines and state-of-the-art methods. We then conduct in-depth ablation studies on key design choices, including the the form of length guidance (Section 5.2) and the statistical metrics for target length selection (Section 5.3). Finally, we provide a mechanistic analysis of how LAPO works, examining its emergent ability for difficulty-aware resource allocation (Section 5.4), its qualitative refinement of reasoning patterns (Section 5.5).
5.1 Main Results
As shown in Table 1, LAPO consistently improves reasoning efficiency across all benchmarks, achieving substantial token reduction while simultaneously enhancing or maintaining accuracy. Our key findings are summarized below.
LAPO simultaneously enhances reasoning performance and reduces test-time computes.
Our framework demonstrates clear superiority over both the foundational models and state-of-the-art baselines. When applied to DeepScaleR-1.5B-Preview, LAPO-I achieves a remarkable 37.8% token reduction while boosting the average accuracy by 2.3 percentage points. This trend holds for DeepSeek-R1-1.5B, where it cuts token usage by 40.9% and improves accuracy by 1.2 points. These results validate that LAPO enables models to generate more concise yet effective reasoning chains.
LAPO surpasses existing efficient reasoning optimization approaches.
Compared to existing methods, LAPO establishes a more effective accuracy-efficiency frontier. On the MATH-500 benchmark, our LAPO-I model (86.3%) substantially outperforms both L1-Max (81.9%) and L1-Exact (80.6%). While L1-Max achieves aggressive token compression, its over-shortening appears to compromise reasoning quality on complex problems. In contrast, LAPO strikes a more effective balance. Against ThinkPrune-4k, which shares our 4k training window, LAPO demonstrates superior performance in both token efficiency and accuracy, confirming its ability to find a better trade-off.
Both Discovery and Internalization stages contribute to the final performance.
The effectiveness of LAPO stems from its two-stage design, where performance progresses verbeter from the Discovery stage (LAPO-D) to the Internalization stage (LAPO-I). LAPO-D first establishes a strong foundation, achieving a significant 36.0% token reduction on its own by learning natural reasoning length distributions. Building on this, LAPO-I achieves an additional 6.9% efficiency gain by internalizing these patterns through in-context guidance. This progressive refinement, combined with LAPO’s consistent performance improvements across diverse benchmarks, confirms that our framework learns generalizable principles of adaptive reasoning rather than task-specific heuristics.
Table 1: Main results on MATH500, AIME2024, AMC, and OlympiadBench. We report Pass@1 accuracy (%) and the average number of generated tokens (#Tok). In each base model group, bold indicates the best and underline indicates the second-best Pass@1 score.
MATH-500 AIME2024 AMC-23 OlympiadBench Average
Pass@1 #Tok Pass@1 #Tok Pass@1 #Tok Pass@1 #Tok Pass@1 #Tok
Base model: DeepSeek-R1-distill-Qwen-1.5B
Base 83.1 4031 30.3 12150 68.3 7222 50.0 8942 57.9 8086
 + LAPO-D 84.7 2566 28.5 8415 72.2 4132 51.3 5595 59.2 5177
  + LAPO-I 84.3 2354 29.3 8318 71.2 3568 51.7 4863 59.1 4775
Base model: DeepScaleR-1.5B-Preview
L1-Exact 80.6 1953 24.4 2625 70.9 2177 48.8 2357 56.2 2278
L1-Max 81.9 1673 24.9 3638 72.7 2705 50.5 2151 57.5 2541
ThinkPrune-I2k 85.5 1707 34.9 5095 74.3 2913 54.7 3498 62.3 3303
ThinkPrune-4k 86.6 2042 35.5 6488 76.3 3839 55.7 4010 63.5 4094
Base 85.8 3280 35.5 9246 74.2 6416 54.6 5974 62.5 6229
 + LAPO-D 86.4 2365 37.6 5945 77.6 3655 56.1 4499 64.4 4116
  + LAPO-I 86.3 2168 38.1 5371 78.3 3765 56.3 4024 64.8 3832
5.2 Ablation Study on the Length Guidance for LAPO-I
To investigate the effectiveness of the in-context length guidance, we conduct an ablation study on the prefill content used in LAPO-I. We compare our default method, Exact Length Guidance (which provides a precise target $n$), against two variants: Range-based Guidance (providing a target range) and Implicit Guidance (no length information, relying only on the reward).
As shown in Table 2, explicit and precise guidance is crucial for performance. Our default method, LAPO-I, achieves the highest average accuracy (64.8%). However, switching to a range-based guidance results in a 0.7% accuracy drop (to 64.1%), and removing the guidance entirely causes a further drop to 63.6%. This experimental results directly validate the effectiveness of the problem-to-length mapping learned in LAPO-D; the more precisely the model is guided to follow these discovered patterns, the better it performs.
Notably, the Implicit Guidance variant not only suffers from lower accuracy but also sees its token consumption revert to a higher 4,088 tokens, nearly identical to the LAPO-D baseline. This demonstrates that the reward signal alone is insufficient to internalize efficient reasoning behaviors. The explicit, in-context guidance acts as a critical bridge, enabling the model to access and leverage the reasoning templates discovered in LAPO-D.
Table 2: Experimental results within different length guidance for LAPO-I. Bold and underline indicate the best and second-best Pass@1 scores. w/ indicates length guidance used in LAPO-I.
Method MATH-500 AIME2024 AMC-23 OlympiadBench Average
Pass@1 #Tok Pass@1 #Tok Pass@1 #Tok Pass@1 #Tok Pass@1 #Tok
Base model: DeepScaleR-1.5B-Preview
Base 85.8 3280 35.5 9246 74.2 6416 54.6 5974 62.5 6229
 LAPO-D 86.4 2365 37.6 5945 77.6 3655 56.1 4499 64.4 4116
  w/ Exact 86.3 2168 38.1 5371 78.3 3765 56.3 4024 64.8 3832
  w/ Range 86.6 2153 36.5 6095 76.9 3600 56.2 4011 64.1 3964
  w/ Implicit 86.9 2181 36.2 5963 76.1 4002 55.1 4206 63.6 4088
Table 3: Experimental results within different statistical metrics used for target length selection in LAPO-I. Bold and underline indicate the best and second-best Pass@1 scores. w/ indicates statistical metrics used for target length selection in LAPO-I.
Method MATH-500 AIME2024 AMC-23 OlympiadBench Average
Pass@1 #Tok Pass@1 #Tok Pass@1 #Tok Pass@1 #Tok Pass@1 #Tok
Base model: DeepScaleR-1.5B-Preview
Base 85.8 3280 35.5 9246 74.2 6416 54.6 5974 62.5 6229
 LAPO-D 86.4 2365 37.6 5945 77.6 3655 56.1 4499 64.4 4116
  w/ Median 86.3 2168 38.1 5371 78.3 3765 56.3 4024 64.8 3832
  w/ Mean 85.6 2308 36.8 6030 77.4 3658 56.6 4164 64.1 4040
  w/ Minimum 85.9 2031 36.3 6080 76.7 3324 55.0 3851 63.5 3821
5.3 Ablation Study on the Statistical Metrics for Target Length Selection
The choice of a statistical measure to derive the target length $n$ from the distribution of successful solutions is critical. We conduct an ablation study comparing three strategies for this selection: using the median (our default), the mean, and the minimum length.
As shown in Table 3, the median proves to be the most effective choice, achieving the best balance between accuracy and efficiency. Using the median as the target yields the highest average accuracy (64.8%) with an efficient token count of 3,832. This validates our hypothesis that the median, due to its robustness to outliers, provides the most representative signal of a “typically” effective reasoning depth. In contrast, the mean is susceptible to a few excessively long, successful solutions, leading it to set overly generous budgets, resulting in higher token usage (4,040) and slightly lower accuracy. The minimum, while achieving the most aggressive compression (3,821 tokens), suffers a significant accuracy drop (to 63.5%), suggesting it promotes an over-shortening strategy that discards necessary reasoning steps. These findings underscore the importance of robust statistical measures for learning a well-calibrated reasoning-efficiency trade-off.
5.4 Difficulty-Aware Computational Allocation
To understand the mechanisms behind LAPO’s efficiency gains, we first examine its macro-level capability to allocate computational resources in proportion to problem complexity. We investigate LAPO’s capacity for difficulty-aware reasoning length allocation by conducting systematic evaluations across mathematical reasoning benchmarks with inherent difficulty gradients: MATH Level 1-5 and AIME 2024. As shown in Figure 4, our experiments reveal a consistent positive correlation between problem complexity and reasoning length across all model variants. This data demonstrates LAPO’s fundamental capacity to proportionally scale computational investment with task demands. Specifically, both LAPO-D and LAPO-I exhibit a near-linear progression in average reasoning length from MATH Level 1 (approximately 1,000-1,200 tokens) to AIME 2024 (5,000-8,500 tokens), validating the proposed LAPO framework’s emergent capability in intrinsic difficulty awareness. The consistent scaling behaviors across different base models (DeepSeek-R1-1.5B and DeepScaleR-1.5B-Pre) and training stages underscore that LAPO develops robust, generalizable heuristics for computational resource management rather than model-specific optimizations.
Figure 3: Reasoning length allocation across mathematical problem difficulty levels. LAPO learns to scale computation with complexity.
Figure 4: Keyword usage of reasoning behaviors across different stages. LAPO selectively prunes hesitant and exploratory thought patterns.
5.5 Qualitative Refinement of Reasoning Behaviors
Beyond quantitatively adjusting how much computation the model uses, we next investigate how it qualitatively refines the reasoning process to achieve this efficiency. To do this, we analyzed the frequency of keywords indicative of different cognitive behaviors, such as “Self-Correction and Verification”, “Exploration and Alternatives”, “Context Setting”, and “Conclusion Drawing”, in the generated responses (Figure 4). The results reveal a significant shift in the model’s reasoning style. The most significant change is a dramatic reduction in keywords associated with Self-Correction and Exploration. The base model (DeepScaleR-1.5B-Pre) frequently employs these terms, indicating a verbose and deliberative reasoning style. After LAPO-D training, and further refined in LAPO-I, the model significantly curtails this internal monologue. This suggests that LAPO effectively discourages redundant verification loops and inefficient exploration of the solution space. Crucially, this efficiency gain is not achieved by indiscriminately shortening the response. The frequency of keywords related to Context Setting and Conclusion Drawing remains stable across training stages. This demonstrates that LAPO selectively prunes inefficient and hesitant thought patterns while preserving the essential scaffolding of a coherent logical argument. The model learns to maintain its ability to properly frame a problem and articulate its deductions.
6 Conclusion
In this work, we introduce Length-Adaptive Policy Optimization (LAPO), a two-stage reinforcement learning framework that enables language models to autonomously adjust reasoning length based on problem complexity. Unlike existing approaches that impose uniform constraints, LAPO recognizes that efficient reasoning requires understanding problem-specific computational needs rather than following rigid rules. Our two-stage design enables a natural progression: models first learn what constitutes appropriate reasoning depth through experience, then develop the ability to anticipate these requirements proactively. This approach mirrors how human experts develop intuition about problem complexity, allocating mental effort proportionally to task demands. Extensive experiments validate LAPO’s effectiveness, achieving remarkable efficiency gains with up to 40.9% reduction in token usage while simultaneously improving accuracy by 2.3% on mathematical reasoning benchmarks. Our analysis reveals that this improvement stems from the model’s ability to distinguish between problems requiring elaborate derivations versus those needing only brief calculations. These results confirm that when models learn from their own successful patterns rather than arbitrary constraints, they develop more robust and efficient reasoning strategies.
Appendix A Technical Appendices and Supplementary Material
A.1 Implementation Details
System prompt used for training. The system prompts used for the two-stage training are shown in the boxes below. The prompt titled LAPO-D-prompt was used for DeepSeek-R1-Distill-Qwen-1.5B, and LAPO-I-prompt was used for DeepScaleR. This approach maintains consistency with the original RL training of DeepSeek-R1.
[…]
Training and Reproduction Details. We trained the model on the OpenRLHF framework. During training, we sampled 8 responses for each query in the batch with a temperature of 1.0, set the kl parameter to 0.0001, used a learning rate of 1e-6 and a batch size of 128, and set the maximum context length to 4K tokens during training. Both LAPO-D and LAPO-I training were conducted for 3 episodes, approximately 240 steps. The $\alpha$ and $\beta$ parameters in $R_{1}$ and $R_{2}$ were 0.7 and 0.8, respectively. All experiments were conducted using 4 A800 GPUs. We provide training hyperparameters in Table 4.
Table 4: Training Hyperparameters
Hyperparameter Value
Epochs 1
Episodes 3
Learning Rate 1e-6
Train Batch Size 128
Temperature 1.0
Rollout per Prompt 8
Prompt Max Length 1024
Generation Max Length 4096
KL Coefficient 0.0001
Precision BF16
$\alpha$ 0.7
$\beta$ 0.8
A.2 Training Dynamics
We analyze the training dynamics by periodically evaluating model checkpoints on the MATH-500 validation set to understand the learning mechanisms of our two-stage framework. As illustrated in Figures 5(a) and 5(b), LAPO achieves a superior balance between efficiency and accuracy across both training stages.
Continuous Efficiency Gains. Figure 5(a) shows a clear, two-step reduction in token generation. In Stage 1, the LAPO-D policy rapidly becomes more concise, with its average length decreasing from a verbose baseline of  3,280 tokens to a stable  2,365 tokens, driven by the length-aware reward ($R_{1}$). Building on this, the LAPO-I policy achieves further compression, reducing the length to below 2,200 tokens. This demonstrates that the plan-adherence reward ($R_{2}$), combined with in-context guidance, effectively encourages the model to execute its self-proposed reasoning plans more precisely.
Accuracy Maintenance and Refinement. Crucially, these efficiency gains do not compromise performance. As shown in Figure 5(b), accuracy on MATH-500 is consistently maintained or improved. The LAPO-D policy’s accuracy climbs from 85.8% to over 86.4%, suggesting the reward mechanism prunes redundant or error-prone reasoning steps. The LAPO-I policy sustains this high accuracy level even on a much tighter token budget. Notably, it exhibits a transient performance peak, a key finding that suggests the in-context guidance actively steers the model toward more focused and effective reasoning, rather than merely acting as a constraint.
In summary, the training dynamics validate our two-stage design. LAPO-D establishes a robust foundation for efficient reasoning, which LAPO-I then refines to achieve a superior performance-cost balance. The smooth convergence on a challenging validation set confirms that by learning from its own successful patterns, the model develops transferable and efficient reasoning strategies.
(a) Average Length on MATH-500
(b) Accuracy on MATH-500
Figure 5: Training dynamics evaluated on the MATH-500 validation set. Checkpoints were saved periodically during training on our mixed dataset. (a) Both LAPO-D and LAPO-I policies learn to significantly reduce the average response length. (b) These efficiency gains are achieved while maintaining or even improving accuracy over the baseline.
A.3 Selection of Training Dataset
As mentioned in section 4 Experiment Setup, we chose a mixed dataset for training in our experiments. In this section, we provide a detailed analysis of the impact of different dataset selections on model performance. Table 5 shows the test results on various benchmarks after two-stage training using different training datasets. Several important findings can be observed from the experimental results. Combined-data achieved the best performance in terms of average accuracy, showing a clear advantage over single-dataset training. This indicates that a dataset with a balanced difficulty distribution helps enhance the model’s generalization ability across different types of questions. In terms of token usage efficiency, the model trained on combined-data also performed the best. This suggests that problems with different difficulty gradients help establish a more accurate complexity-length mapping relationship. By exposing the model to a wider range of problem difficulties, it can better learn the optimal thinking range for different questions. Taking all these factors into consideration, we selected the mixed dataset as the training data to expose the model to a more diverse set of problems and enable it to deeply learn the optimal reasoning patterns for different questions.
Table 5: Ablation study on the training dataset. This table compares performance when trained on different data sources. For each metric column, bold indicates the best score and underline indicates the second-best score across all configurations.
Method MATH500 AIME2024 AMC-23 OlympiadBench Average
Pass@1 #Tok Pass@1 #Tok Pass@1 #Tok Pass@1 #Tok Pass@1 #Tok
Training Data: Combined (Ours)
 LAPO-D 86.4 2365 37.6 5945 77.6 3655 56.1 4499 64.4 4116
 LAPO-I 86.3 2168 38.1 5371 78.3 3765 56.3 4024 64.8 3832
Training Data: DeepScaleR-only
 LAPO-D 86.1 2397 36.8 6153 76.8 3983 55.5 4258 63.8 4197
 LAPO-I 86.1 2210 36.5 6418 77.0 3791 55.6 3933 63.8 4088
Training Data: MATH-only
 LAPO-D 86.5 2398 38.0 7034 77.3 4060 55.8 4494 64.4 4496
 LAPO-I 86.1 2340 35.5 6452 75.8 4021 54.5 4194 63.0 4251

### v2

@@
 Existing approaches to address this challenge fall into three main categories, each with inherent limitations.
-Direct length reduction methods either rely on reward design (Yang et al., 2025; Huang et al., 2025) that can cause over-shortening and accuracy degradation, or impose hard length constraints (Aggarwal & Welleck, 2025; Hou et al., 2025) that lack adaptability across problem types.
+Direct length reduction methods either rely on reward design (Yang et al., 2025; Huang et al., 2025a) that can cause over-shortening and accuracy degradation, or impose hard length constraints (Aggarwal & Welleck, 2025; Hou et al., 2025) that lack adaptability across problem types.
 Dynamic early-stopping approaches (Qiao et al., 2025; Muennighoff et al., 2025) make real-time termination decisions but often truncate mid-reasoning, disrupting the thinking process.
-Adaptive thinking methods (Lou et al., 2025; Zhang et al., 2025; Fang et al., 2025) enable models to switch between thinking and non-thinking modes but operate at a coarse granularity.
+Adaptive thinking methods (Lou et al., 2025; Zhang et al., 2025a; Fang et al., 2025a) enable models to switch between thinking and non-thinking modes but operate at a coarse granularity.
 The fundamental limitation of these approaches lies in their treatment of length control as an external constraint imposed upon the reasoning process.
@@
 By incorporating target length information directly into the problem prompt, we enable models to plan their reasoning trajectory before beginning the solution process.
-Crucially, to ensure models can reason adaptively without requiring predefined lengths at inference time, we embed the length constraint as a self-declarative statement immediately after the ¡think¿ token.
+Crucially, to ensure models can reason adaptively without requiring predefined lengths at inference time, we embed the length constraint as a self-declarative statement immediately after the <think> token.
 This technique reframes the budget not as an external command, but as part of the model’s own internal reasoning plan.
@@
 We derive two key statistics from these rollouts. First, we establish a reasonable length range using percentiles to filter outliers while preserving central tendencies:
-$[L_{\min},L_{\max}]=[\text{Percentile}_{30}(\mathcal{L}_{q}),\text{Percentile}_{70}(\mathcal{L}_{q})]$ (2)
+$L_{\min},L_{\max}]=[\text{Percentile}_{30}(\mathcal{L}_{q}),\text{Percentile}_{70}(\mathcal{L}_{q})]$ (2)
 Second, we create a problem-to-length mapping that will guide the Internalization stage:
@@
 We employ a composite reward function balancing accuracy and efficiency:
-$R_{1}(r_{i},q)=\mathbb{I}(y_{i}=y_{\text{gold}})+\alpha\cdot R_{\text{length-1}}(r_{i},q)$ (4)
-The length component operates on a crucial principle—only correct responses receive length-based rewards:
-$R_{\text{length-1}}(r_{i},q)=\begin{cases}1.0&\text{if correct and }L_{\min}\leq|r_{i}|\leq L_{\max},\\ \max(0,1-d_{i}/100)&\text{if correct and }|r_{i}|\notin[L_{\min},L_{\max}],\\ 0&\text{if incorrect;}\end{cases}$ (5)
-where $d_{i}=\min(||r_{i}|-L_{\min}|,||r_{i}|-L_{\max}|)$.
+$R_{D}(r_{i},q)=\mathbb{I}(y_{i}=y_{\text{gold}})+\alpha\cdot R_{\text{1}}(r_{i},q)$ (4)
+The length component operates on a crucial principle—only correct responses receive length-based rewards.
+Let $\mathcal{C}_{i}=\mathbb{I}(y_{i}=y_{\text{gold}})$ indicate whether the response is correct, and define the distance to the target length range as $d_{i}=\min(||r_{i}|-L_{\min}|,||r_{i}|-L_{\max}|)$.
+We introduce a linear decay function $f(d)=\max(0,1-d/100)$ to penalize deviations from the efficient length range.
+The length reward is then defined as:
+$R_{\text{1}}(r_{i},q)=\begin{cases}1.0&\text{if }\mathcal{C}_{i}=1\land|r_{i}|\in[L_{\min},L_{\max}]\\ f(d_{i})&\text{if }\mathcal{C}_{i}=1\land|r_{i}|\notin[L_{\min},L_{\max}]\\ 0&\text{if }\mathcal{C}_{i}=0\end{cases}$ (5)
 This design creates gradients guiding models toward efficient lengths while allowing flexibility for complex problems.
@@
 We augment each problem prompt with explicit length guidance:
-$\text{prompt}_{q}^{\prime}=\text{prompt}_{q}+\text{``{<think>} I will answer the question with }n\text{ tokens.''}$.” (6)
+$_{q}^{\prime} = prompt$q + “<think> I will answer the question with $n$ tokens.”
 where $n=\mathcal{M}(q)$ from the Discovery stage.
@@
 Length-Adherence Reward.
-To encourage the model to follow its self-declared reasoning budget from Equation 6, the Internalization stage employs a precision-focused reward function.
+To encourage the model to follow its self-declared reasoning budget, the Internalization stage employs a precision-focused reward function.
 This function is designed to reward the alignment between the model’s output length and its self-declared budget n.
 The total reward is defined as:
-$R_{2}(r_{i},q^{\prime})=\mathbb{I}(y_{i}=y_{\text{gold}})+\beta\cdot R_{\text{Length-2}}(r_{i},q^{\prime})$ (7)
-where the adherence component, $R_{\text{Length-2}}$, is only granted for correct solutions:
-$R_{\text{Length-2}}(r_{i},n)=\begin{cases}\exp\left(-\frac{(|r_{i}|-n)^{2}}{2\sigma^{2}}\right)&\text{if correct},\\ 0&\text{otherwise};\end{cases}$ (8)
+$R_{I}(r_{i},q^{\prime})=\mathbb{I}(y_{i}=y_{\text{gold}})+\beta\cdot R_{\text{2}}(r_{i},q^{\prime})$ (6)
+where the adherence component, $R_{\text{2}}$, is only granted for correct solutions:
+$R_{\text{2}}(r_{i},n)=\begin{cases}\exp\left(-\frac{(|r_{i}|-n)^{2}}{2\sigma^{2}}\right)&\text{if }\mathcal{C}_{i}=1,\\ 0&\text{if }\mathcal{C}_{i}=0;\end{cases}$ (7)
 This Gaussian-inspired reward positively reinforces solutions that are both correct and consistent with the intended reasoning depth.
@@
 7:  for each problem $q\in\mathcal{B}$ do
-8:         Generate $N$ rollouts: $\{r_{1},...,r_{N}\}\sim\pi_{\theta}(q)$
+8:   Generate NN rollouts: $\{r_{1},\dots,r_{N}\}\sim\pi_{\theta}(q)$
 9:   Collect correct lengths: $\mathcal{L}_{q}=\{|r_{i}|:y_{i}=y_{\text{gold}}\}$
-10:         Compute range: $[L_{\min},L_{\max}]=[\text{P}_{30}(\mathcal{L}_{q}),\text{P}_{70}(\mathcal{L}_{q})]$
+10:   Compute range: $L_{\min},L_{\max}]=[\text{P}_{30}(\mathcal{L}_{q}),\text{P}_{70}(\mathcal{L}_{q})]$
 11:   Update mapping: $\mathcal{M}(q)=\text{Median}(\mathcal{L}_{q})$
-12:         Compute rewards: $R_{1}(r_{i},q)=\mathbb{I}(y_{i}=y_{\text{gold}})+\alpha\cdot R_{\text{length-1}}(r_{i},q)$
+12:   $\text{Compute rewards: }R_{D}(r_{i},q)={}\mathbb{I}(y_{i}=y_{\text{gold}})+\alpha\cdot R_{\text{1}}(r_{i},q)$
 13:  end for
@@
 20:  for each problem $q\in\mathcal{B}$ do
-21:         Augment prompt: $q^{\prime}=q+\text{``{<think>} I will answer the question with }\mathcal{M}(q)\text{ tokens.''}$.”
-22:         Generate $N$ rollouts: $\{r_{1},...,r_{N}\}\sim\pi_{\theta}(q^{\prime})$
-23:         Compute rewards: $R_{2}(r_{i},q^{\prime})=\mathbb{I}(y_{i}=y_{\text{gold}})+\beta\cdot R_{\text{Length-2}}(r_{i},q^{\prime})$
+21:   Augment prompt: $q^{\prime}\leftarrow{}q+\text{``{<think>} }\text{I will}$  $\text{ answer the question}\text{ with }\mathcal{M}(q)\text{ tokens.''}$
+22:   Generate NN rollouts: $\{r_{1},\dots,r_{N}\}\sim\pi_{\theta}(q^{\prime})$
+23:   $\text{Compute rewards: }R_{I}(r_{i},q^{\prime})={}\mathbb{I}(y_{i}=y_{\text{gold}})+\beta\cdot R_{\text{2}}(r_{i},q^{\prime})$
 24:    Update mapping $\mathcal{M}(q)$ using dual-strategy (Eq. 8)
@@
 A cornerstone of our framework is how it fosters genuine internalization, enabling inference-time flexibility without explicit length targets.
-The key lies in the design of the augmented prompt (Equation 6).
-By placing the length guidance ”I will answer the question with $n$ tokens.” immediately after the <think> token, we transform an external constraint into an intrinsic part of the model’s cognitive process.
+The key lies in the design of the augmented prompt.
+Placing the self-declarative guidance immediately after the <think> token transforms an external constraint into an intrinsic part of the model’s cognitive plan.
 During the Internalization stage, we refine $\mathcal{M}$ based on new GRPO rollouts with a dual-strategy update:
@@
 Discovery Stage
-(Lines 4-13): The model explores natural reasoning patterns through GRPO training with length-aware rewards.
+(Lines 4-15): The model explores natural reasoning patterns through GRPO training with length-aware rewards.
 For each problem in the training batch, we generate multiple rollouts and extract statistics from successful responses.
@@
 Internalization Stage
-(Lines 15-24): The model learns to internalize efficient reasoning by incorporating discovered length patterns into the training process.
+(Lines 17-27): The model learns to internalize efficient reasoning by incorporating discovered length patterns into the training process.
 Each problem prompt is augmented with target length information derived from the Discovery stage.
@@
 4 Experiment Setup
+5 Experiment Setup
 Training Details.
-We train our models on a mixed dataset of 10,000 mathematical problems.
-This dataset comprises 6,000 examples randomly sampled from the DeepScaleR-Preview-Dataset (Luo et al., 2025b), which sources problems from competitive math contests (e.g., AIME, AMC) and benchmarks (e.g., Omni-Math (Gao et al., 2024), STILL (Min et al., 2024)).
-The remaining 4,000 examples are drawn from the intermediate levels (3-5) of the MATH dataset (Hendrycks et al., 2020) to ensure a balanced difficulty distribution.
-For the base models, we experiment with DeepSeek-R1-1.5B(Guo et al., 2025), a strong instruction-tuned reasoner, and DeepScaleR-1.5B-Preview(Luo et al., 2025b), an advanced RL-finetuned version known for high-quality but verbose outputs.
+We train our models on a mixed dataset of 10,000 mathematical problems to ensure a balanced difficulty distribution, comprising 6,000 examples from the DeepScaleR-Preview-Dataset Luo et al.
+(2025b) and 4,000 from the intermediate levels of the MATH dataset Hendrycks et al.
+(2020).
+We apply LAPO to two base models: DeepSeek-R1-1.5B Guo et al.
+(2025) and DeepScaleR-1.5B-Preview Luo et al.
+(2025b).
 All models are trained using the Group Relative Policy Optimization (GRPO) algorithm.
-The LAPO framework is applied in two sequential stages.
-Stage 1 (LAPO-Discovery) is trained for 240 steps (approx.
-3 epochs) with the length-aware reward $R_{1}$ (Eq.
+Discovery Stage(LAPO-D) is trained for 3 episodes with reward $R_{D}$ (Eq.
 4), where hyperparameter $\alpha$ is 0.7.
-The resulting model then serves as the initialization for Stage 2 (LAPO-Internalization), which is trained for another 240 steps using the plan-adherence reward $R_{2}$ (Eq.
-7) with $\beta$ set to 0.8.
-Due to computational constraints, the maximum context length is limited to 4,096 tokens during training.
-More training details are provided in Appendix A.1.
+The resulting model then serves as the initialization for the subsequent Internalization Stage(LAPO-I), which is trained for 3 episodes using reward $R_{I}$ (Eq.
+6) with $\beta$ set to 0.7.
+These hyperparameters were empirically chosen to provide a strong incentive for efficiency while ensuring correctness remains the primary learning signal.
+Note that we did not conduct extensive hyperparameter tuning, so one can expect further improvements with additional optimization.
+Besides, due to computational constraints, the maximum context length is limited to 4,096 tokens during training.
+Crucially, to ensure a fair comparison, most baselines, including ThinkPrune and L1, were also trained or evaluated under this same 4k context limit.
+Further hyperparameters are detailed in Appendix.
 Evaluation Details.
-We evaluate the trained models on four challenging benchmarks representing a spectrum of mathematical difficulty: MATH-500 (Hendrycks et al., 2020), AIME2024, AMC23, and Olympiad-Bench (He et al., 2024).
-Following standard practices (Guo et al., 2025), we use a sampling temperature of 0.6, a top-p of 0.95, and a maximum generation length of 32,768 tokens.
-For each benchmark, we report Pass@1 accuracy and the average number of tokens per response.
-The number of samples per query is 4 for MATH-500 and Olympiad-Bench, and 32 for the more competitive AIME2024 and AMC23.
+At inference, we expand the generation window to a generous 32,768 tokens for all models to assess their true, unconstrained reasoning capabilities.
+This setup allows us to isolate the efficiency gains stemming directly from the LAPO framework, rather than from simple context window limitations.
+We evaluate on four challenging benchmarks: MATH-500 Hendrycks et al.
+(2020), AIME2024, AMC23, and Olympiad-Bench He et al.
+(2024).
+Following standard practices Guo et al.
+(2025), we report both Pass@1 accuracy and the average number of tokens.
+For each problem, we sample N responses (4 for MATH-500/OlympiadBench, 32 for AIME/AMC) with a temperature of 0.6 and a top-p of 0.95.
 Baselines.
-We benchmark LAPO against two classes of baselines: the foundational models we build upon and existing methods designed for efficient reasoning.
+We benchmark LAPO against three classes of baselines: the foundational models, an ablation baseline, and existing methods designed for efficient reasoning.
 First, we evaluate the Base Models to establish a performance starting point.
-These include DeepSeek-R1-1.5B(Guo et al., 2025), a strong instruction-tuned model, and DeepScaleR-1.5B-Preview(Luo et al., 2025b), a powerful RL-finetuned reasoner known for its high-quality but often verbose outputs.
-Both are evaluated without any length-control modifications.
-Second, we compare against state-of-the-art Efficient Reasoning Baselines that represent two distinct technical approaches.
-For methods focusing on direct length budgeting, we evaluate L1-Exact and L1-Max(Aggarwal & Welleck, 2025).
-Both train a model to adhere to a user-specified length constraint by penalizing deviations from the target budget, with L1-Exact enforcing a precise match and L1-Max treating the budget as a soft upper limit.
-For the iterative pruning strategy, we compare against ThinkPrune(Hou et al., 2025).
-This method trains a model to shorten its reasoning by imposing a hard token limit during its iterative RL process, where any response exceeding the limit receives a zero reward.
-We evaluate its two publicly available variants, ThinkPrune-iter2k and ThinkPrune-4k.
+These include DeepSeek-R1-1.5B Guo et al.
+(2025) and DeepScaleR-1.5B-Preview Luo et al.
+(2025b).
+Second, to isolate the effect of our length-reward, we also include an Ablation Baseline, denoted as Acc-Only, which is trained with GRPO using only the accuracy reward.
+Finally, we compare against several state-of-the-art Efficient Reasoning Baselines, which represent different philosophies for achieving efficiency.
+(1)Implicit Regularization: HAPO Huang et al.
+(2025b), which uses history-aware rewards.
+(2)Budget-Driven Control: L1 Aggarwal & Welleck (2025) and ThinkPrune Hou et al.
+(2025), which follow external length targets.
+(3)Adaptive Activation: AutoThink Tu et al.
+(2025), AdaptThink Zhang et al.
+(2025b), and Thinkless Fang et al.
+(2025b), which learn a binary think/no-think policy.
 6 Results and Analysis
-We present comprehensive experimental results and analysis to validate LAPO’s effectiveness and understand its underlying mechanisms.
-We begin with the main results (Section 5.1), benchmarking LAPO against baselines and state-of-the-art methods.
-We then conduct in-depth ablation studies on key design choices, including the the form of length guidance (Section 5.2) and the statistical metrics for target length selection (Section 5.3).
-Finally, we provide a mechanistic analysis of how LAPO works, examining its emergent ability for difficulty-aware resource allocation (Section 5.4), its qualitative refinement of reasoning patterns (Section 5.5).
-5.1 Main Results
-As shown in Table 1, LAPO consistently improves reasoning efficiency across all benchmarks, achieving substantial token reduction while simultaneously enhancing or maintaining accuracy.
-Our key findings are summarized below.
-LAPO simultaneously enhances reasoning performance and reduces test-time computes.
-Our framework demonstrates clear superiority over both the foundational models and state-of-the-art baselines.
-When applied to DeepScaleR-1.5B-Preview, LAPO-I achieves a remarkable 37.8% token reduction while boosting the average accuracy by 2.3 percentage points.
-This trend holds for DeepSeek-R1-1.5B, where it cuts token usage by 40.9% and improves accuracy by 1.2 points.
-These results validate that LAPO enables models to generate more concise yet effective reasoning chains.
-LAPO surpasses existing efficient reasoning optimization approaches.
-Compared to existing methods, LAPO establishes a more effective accuracy-efficiency frontier.
-On the MATH-500 benchmark, our LAPO-I model (86.3%) substantially outperforms both L1-Max (81.9%) and L1-Exact (80.6%).
-While L1-Max achieves aggressive token compression, its over-shortening appears to compromise reasoning quality on complex problems.
-In contrast, LAPO strikes a more effective balance.
-Against ThinkPrune-4k, which shares our 4k training window, LAPO demonstrates superior performance in both token efficiency and accuracy, confirming its ability to find a better trade-off.
-Both Discovery and Internalization stages contribute to the final performance.
-The effectiveness of LAPO stems from its two-stage design, where performance progresses verbeter from the Discovery stage (LAPO-D) to the Internalization stage (LAPO-I).
-LAPO-D first establishes a strong foundation, achieving a significant 36.0% token reduction on its own by learning natural reasoning length distributions.
-Building on this, LAPO-I achieves an additional 6.9% efficiency gain by internalizing these patterns through in-context guidance.
-This progressive refinement, combined with LAPO’s consistent performance improvements across diverse benchmarks, confirms that our framework learns generalizable principles of adaptive reasoning rather than task-specific heuristics.
-Table 1: Main results on MATH500, AIME2024, AMC, and OlympiadBench.
+Table 1: Main results on MATH500, AIME2024, AMC23, and OlympiadBench.
 We report Pass@1 accuracy (%) and the average number of generated tokens (#Tok).
-In each base model group, bold indicates the best and underline indicates the second-best Pass@1 score.
+For each metric, bold indicates the best and underline indicates the second-best Pass@1 score within each base model group.
 MATH-500 AIME2024 AMC-23 OlympiadBench Average
 Pass@1 #Tok Pass@1 #Tok Pass@1 #Tok Pass@1 #Tok Pass@1 #Tok
-Base model: DeepSeek-R1-distill-Qwen-1.5B
+Base model: DeepSeek-R1-1.5B
+HAPO 82.2 2288 31.3 8649 67.3 4735 50.1 5024 57.7 5174
+AutoThink 83.5 2017 29.7 7084 70.2 3499 51.2 4606 58.6 3825
+AdaptThink 81.6 1580 23.9 6432 63.2 2860 48.5 4616 54.3 3871
 Base 83.1 4031 30.3 12150 68.3 7222 50.0 8942 57.9 8086
+ + Acc-Only 83.3 3061 31.6 10628 70.5 5307 50.6 6402 59.0 6349
  + LAPO-D 84.7 2566 28.5 8415 72.2 4132 51.3 5595 59.2 5177
@@
 ThinkPrune-4k 86.6 2042 35.5 6488 76.3 3839 55.7 4010 63.5 4094
+HAPO 84.4 2370 31.4 7702 70.3 4301 51.4 4571 59.3 4736
+AutoThink 84.9 1635 36.2 7201 67.8 3658 52.5 4085 60.4 4144
+Thinkless 81.3 2944 28.9 9143 65.7 5276 50.2 6057 56.5 5855
 Base 85.8 3280 35.5 9246 74.2 6416 54.6 5974 62.5 6229
+ + Acc-Only 85.6 2510 36.9 7319 77.6 4244 55.6 4712 63.9 4696
  + LAPO-D 86.4 2365 37.6 5945 77.6 3655 56.1 4499 64.4 4116
   + LAPO-I 86.3 2168 38.1 5371 78.3 3765 56.3 4024 64.8 3832
-5.2 Ablation Study on the Length Guidance for LAPO-I
-To investigate the effectiveness of the in-context length guidance, we conduct an ablation study on the prefill content used in LAPO-I.
-We compare our default method, Exact Length Guidance (which provides a precise target $n$), against two variants: Range-based Guidance (providing a target range) and Implicit Guidance (no length information, relying only on the reward).
-As shown in Table 2, explicit and precise guidance is crucial for performance.
-Our default method, LAPO-I, achieves the highest average accuracy (64.8%).
-However, switching to a range-based guidance results in a 0.7% accuracy drop (to 64.1%), and removing the guidance entirely causes a further drop to 63.6%.
-This experimental results directly validate the effectiveness of the problem-to-length mapping learned in LAPO-D; the more precisely the model is guided to follow these discovered patterns, the better it performs.
-Notably, the Implicit Guidance variant not only suffers from lower accuracy but also sees its token consumption revert to a higher 4,088 tokens, nearly identical to the LAPO-D baseline.
-This demonstrates that the reward signal alone is insufficient to internalize efficient reasoning behaviors.
-The explicit, in-context guidance acts as a critical bridge, enabling the model to access and leverage the reasoning templates discovered in LAPO-D.
-Table 2: Experimental results within different length guidance for LAPO-I. Bold and underline indicate the best and second-best Pass@1 scores. w/ indicates length guidance used in LAPO-I.
+We present comprehensive experimental results and analysis to validate LAPO’s effectiveness and understand its underlying mechanisms.
+We begin with the main results (Section 6.1), benchmarking LAPO against baselines and state-of-the-art methods.
+We then conduct in-depth ablation studies on key design choices, including the the form of length guidance (Section 6.2) and the statistical metrics for target length selection (Section 6.3).
+And a targeted experiment demonstrating the model’s robust internalization of reasoning efficiency (Section 6.4).
+Finally, we provide a mechanistic analysis of how LAPO works, examining its emergent ability for difficulty-aware resource allocation (Section 6.5), its qualitative refinement of reasoning patterns (Section 6.6).
+6.1 Main Results
+As shown in Table 1, LAPO achieves a superior balance of reasoning accuracy and computational efficiency, consistently outperforming its base models and establishing a new state-of-the-art frontier among methods that do not rely on external length controls.
+LAPO simultaneously enhances reasoning performance and reduces test-time computes.
+Compared to its base models, LAPO delivers substantial gains.
+On DeepScaleR-1.5B-Preview, it reduces tokens by 38.5% while boosting average accuracy by 2.3 points; a similar trend holds for DeepSeek-R1-1.5B (41.0% token cut and 1.2 point accuracy gain).
+This validates that LAPO learns to produce more concise yet effective reasoning.
+LAPO surpasses existing efficient reasoning optimization approaches.
+When compared with leading efficiency methods, LAPO consistently demonstrates a superior accuracy-efficiency trade-off.
+On the more capable DeepScaleR-1.5B base model, LAPO-I achieves the highest average accuracy among all tested methods.
+This advantage holds across different baseline paradigms.
+It surpasses budget-driven methods like ThinkPrune-4k and L1-Max under a fair 4k training context.
+Compared to implicit regularization methods like HAPO, LAPO shows a clear advantage in preserving accuracy.
+Furthermore, while adaptive activation methods like AutoThink can be highly token-efficient, they do not reach the same level of reasoning quality.
+This comprehensive comparison highlights that LAPO’s “Discover-Internalize” process, which fosters an autonomous and continuous length adaptation, leads to a more robust and effective reasoning policy than methods relying on external budgets, progressive compression, or binary mode-switching.
+Both Discovery and Internalization stages contribute to the final performance.
+LAPO-D first establishes a strong foundation, achieving a significant 36.0% token reduction on its own by learning natural reasoning length distributions.
+This is highlighted by comparing it to the Acc-Only baseline.
+While simply finetuning for accuracy yields some token reduction, LAPO-D’s length-aware reward achieves substantially greater efficiency while also improving average accuracy by 0.5 points.
+This demonstrates that encouraging conciseness via our reward not only prunes redundant thoughts but also helps the model find more robust reasoning patterns.
+Building on this superior foundation, LAPO-I achieves an additional 6.9% efficiency gain by internalizing these patterns through in-context guidance.
+This progressive refinement indicates that our framework learns a generalizable principle of adaptive reasoning.
+6.2 Ablation Study on In-Context Guidance
+To validate that our method’s success stems from internalizing a self-proposed plan, we ablate the two key factors of our in-context guidance: its form (how precise the guidance is) and its position (whether it’s part of the model’s internal thought process).
+We compare our default approach (w/ Exact) against three variants: w/ Range (less precise guidance), w/ Outside (placing the guidance before <think>), and w/ Implicit (no guidance, relying only on the reward).
+As shown in Table 2, the results demonstrate that both form and position are critical for effective internalization.
+Our default method outperforms the less precise Range variant, indicating that specific targets discovered in Discovery stage provide a stronger learning signal.
+More critically, the guidance’s position determines whether the model internalizes a plan or merely follows instructions.
+Moving the guidance outside the <think> block transforms it into an external command and causes accuracy to drop significantly to 63.9%.
+This illustrates that the model performs best when the budget is framed as part of its own cognitive plan.
+Finally, removing the guidance entirely results in the worst performance, with accuracy dropping to 63.6% and token count reverting to the LAPO-D baseline.
+This indicates that our explicit, properly-positioned, self-declarative guidance is the critical mechanism for internalization.
+Table 2: Experimental results with different length guidance for LAPO-I. Bold and underline indicate the best and second-best Pass@1 scores. w/ indicates the length guidance used in LAPO-I.
 Method MATH-500 AIME2024 AMC-23 OlympiadBench Average
@@
   w/ Range 86.6 2153 36.5 6095 76.9 3600 56.2 4011 64.1 3964
+  w/ Outside 86.5 2251 36.4 5882 76.3 3850 55.4 4105 63.9 4022
   w/ Implicit 86.9 2181 36.2 5963 76.1 4002 55.1 4206 63.6 4088
@@
   w/ Minimum 85.9 2031 36.3 6080 76.7 3324 55.0 3851 63.5 3821
-5.3 Ablation Study on the Statistical Metrics for Target Length Selection
+6.3 Ablation on Statistical Metrics for Target Length
 The choice of a statistical measure to derive the target length $n$ from the distribution of successful solutions is critical.
@@
 These findings underscore the importance of robust statistical measures for learning a well-calibrated reasoning-efficiency trade-off.
+Table 4: Robustness of LAPO-I to conflicting length instructions on MATH-500.
+Method Length Constraint MATH-500
+Pass@1 (%) #Tok
+LAPO-I
+Base N/A 86.3 2168
+   +Short 500 86.0 2279
+   +Long 3500 85.9 2300
+LAPO-I w/ Outside
+Base N/A 86.2 2251
+   +Short 500 85.1 1247
+   +Long 3500 86.1 2821
+6.4 Analysis of Internalization
+To validate that LAPO fosters genuine internalization, we stress-tested our default LAPO-I model against the w/ Outside ablation variant using adversarial Short (500 tokens) and Long (3500 tokens) length prompts.
+The results in Table 4 reveal a stark behavioral divergence.
+Our default LAPO-I remains robust, its output length staying stable around its 2200-token baseline, thus ignoring the conflicting external instructions.
+In contrast, the w/ Outside model is clearly influenced: its token count drops to 1247 under the Short constraint and rises to 2821 under the Long one.
+This comparison indicates that the placement of guidance is critical.
+Framing the budget as part of the model’s internal plan (inside <think>) builds a robust, internalized behavior.
+Framing it externally teaches superficial instruction-following.
+This indicates the observed robustness of LAPO-I is a direct result of our internalization mechanism.
 6.5 Difficulty-Aware Computational Allocation
-To understand the mechanisms behind LAPO’s efficiency gains, we first examine its macro-level capability to allocate computational resources in proportion to problem complexity.
-We investigate LAPO’s capacity for difficulty-aware reasoning length allocation by conducting systematic evaluations across mathematical reasoning benchmarks with inherent difficulty gradients: MATH Level 1-5 and AIME 2024.
-As shown in Figure 4, our experiments reveal a consistent positive correlation between problem complexity and reasoning length across all model variants.
-This data demonstrates LAPO’s fundamental capacity to proportionally scale computational investment with task demands.
-Specifically, both LAPO-D and LAPO-I exhibit a near-linear progression in average reasoning length from MATH Level 1 (approximately 1,000-1,200 tokens) to AIME 2024 (5,000-8,500 tokens), validating the proposed LAPO framework’s emergent capability in intrinsic difficulty awareness.
-The consistent scaling behaviors across different base models (DeepSeek-R1-1.5B and DeepScaleR-1.5B-Pre) and training stages underscore that LAPO develops robust, generalizable heuristics for computational resource management rather than model-specific optimizations.
+To understand the mechanisms behind LAPO’s efficiency gains, we examine its ability to allocate computational resources in proportion to problem complexity.
+We evaluate LAPO-trained models on benchmarks with clear difficulty gradients, from MATH Level 1 up to the highly complex AIME 2024.
+As shown in Figure 4, our models demonstrate a remarkable emergent capability for difficulty-aware resource allocation.
+There is a clear, near-linear positive correlation between problem complexity and the average reasoning length.
+On simpler problems, the models generate concise responses, while for the most challenging AIME questions, they produce extensive reasoning chains that are substantially longer than any solution observed during the training phase.
+This ability to extrapolate reasoning depth well beyond the bounds of their training experience is a crucial finding.
+It provides strong evidence that LAPO does not merely teach models to compress their outputs.
+Instead, it successfully imparts a generalizable principle of complexity-to-length mapping.
+This allows the models to dynamically and appropriately scale their computational investment when faced with novel problems of varying difficulty.
+The consistent scaling behavior across different base models further underscores that LAPO develops a robust, fundamental reasoning strategy rather than model-specific optimizations.
 Figure 3: Reasoning length allocation across mathematical problem difficulty levels. LAPO learns to scale computation with complexity.
@@
 $\alpha$ 0.7
-$\beta$ 0.8
+$\beta$ 0.7
 A.2 Training Dynamics
@@
 (b) These efficiency gains are achieved while maintaining or even improving accuracy over the baseline.
-A.3 Selection of Training Dataset
-As mentioned in section 4 Experiment Setup, we chose a mixed dataset for training in our experiments.
-In this section, we provide a detailed analysis of the impact of different dataset selections on model performance.
-Table 5 shows the test results on various benchmarks after two-stage training using different training datasets.
-Several important findings can be observed from the experimental results.
-Combined-data achieved the best performance in terms of average accuracy, showing a clear advantage over single-dataset training.
-This indicates that a dataset with a balanced difficulty distribution helps enhance the model’s generalization ability across different types of questions.
-In terms of token usage efficiency, the model trained on combined-data also performed the best.
-This suggests that problems with different difficulty gradients help establish a more accurate complexity-length mapping relationship.
-By exposing the model to a wider range of problem difficulties, it can better learn the optimal thinking range for different questions.
-Taking all these factors into consideration, we selected the mixed dataset as the training data to expose the model to a more diverse set of problems and enable it to deeply learn the optimal reasoning patterns for different questions.
 Table 6: Ablation study on the training dataset.
@@
  LAPO-I 86.1 2340 35.5 6452 75.8 4021 54.5 4194 63.0 4251
+A.3 Selection of Training Dataset
+As mentioned in section 4 Experiment Setup, we chose a mixed dataset for training in our experiments.
+In this section, we provide a detailed analysis of the impact of different dataset selections on model performance.
+Table 6 shows the test results on various benchmarks after two-stage training using different training datasets.
+Several important findings can be observed from the experimental results.
+Combined-data achieved the best performance in terms of average accuracy, showing a clear advantage over single-dataset training.
+This indicates that a dataset with a balanced difficulty distribution helps enhance the model’s generalization ability across different types of questions.
+In terms of token usage efficiency, the model trained on combined-data also performed the best.
+This suggests that problems with different difficulty gradients help establish a more accurate complexity-length mapping relationship.
+By exposing the model to a wider range of problem difficulties, it can better learn the optimal thinking range for different questions.
+Taking all these factors into consideration, we selected the mixed dataset as the training data to expose the model to a more diverse set of problems and enable it to deeply learn the optimal reasoning patterns for different questions.
+Table 7: Performance on the GPQA benchmark. LAPO demonstrates generalizable efficiency and accuracy gains in a non-mathematical, knowledge-intensive domain.
+Method Pass@1 (%) #Tokens
+Base Model: DeepSeek-R1-1.5B
+Base 36.1 10297
+   + LAPO-D 38.1 7596
+   + LAPO-I 36.9 7235
+Base Model: DeepScaleR-1.5B-Preview
+Base 36.1 7667
+   + LAPO-D 38.3 6176
+   + LAPO-I 37.8 6154
+A.4 Generalizability to Expert-Level Question Answering.
+To test if LAPO’s benefits extend beyond structured mathematical reasoning, we evaluated our method on the GPQA benchmark.
+The results, presented in Table 7, demonstrate that LAPO’s core principles are highly generalizable.
+For both base models, LAPO achieves a compelling dual improvement in accuracy and efficiency.
+On the DeepSeek-R1-1.5B model, LAPO-D improves Pass@1 accuracy by a significant 2.0 points while reducing token generation by 26.2%.
+Similarly, on the more advanced DeepScaleR-1.5B-Preview, LAPO-D boosts accuracy by 2.2 points and cuts tokens by 19.4%.
+The internalization stage consistently pushes efficiency further while maintaining a strong accuracy improvement over the baseline.
+This robust performance on a knowledge-intensive, non-mathematical task indicates that LAPO is not merely exploiting domain-specific patterns.
+Instead, it learns a fundamental and transferable skill: how to allocate cognitive effort efficiently for complex reasoning across different domains.